3x^3+x^2-3x-1=0 (The ^3 and ^2 and exponents)

I have tried to factor and pull out the x but I am getting confused with the two different power. I got x=-1, x=-1/3 and x=1 but i dont think i need three answers

Answer by stanbon(75887)

3x^3+x^2-3x-1=0
x^2(x+1) -3(x+1) = 0
(x+1)(x^2-3) = 0
x = -1 or x=sqrt(3) or x=-sqrt(3)
=====================================
Cheers,
Stan H.